OpenStudy (anonymous):

Use the properties of logarithms to write the following as sum and/or difference of logarithms: log√[(x^2 -1)/(x^3)((y^2+1)^5)]

OpenStudy (anonymous):

ok i guess we can straighten out this mess

OpenStudy (anonymous):

is the square root over the whole input, or just in the numerator

OpenStudy (anonymous):

It's over the whole thing

OpenStudy (anonymous):

the the first step is to get rid of the square root by putting a \(\frac{1}{2}\) out front

OpenStudy (anonymous):

\[\frac{1}{2}\log\left(\frac{(x^2-1)}{x^3(y^2+1)^5}\right)\]

OpenStudy (anonymous):

the turn the division in to a subtraction \[\frac{1}{2}\log(x^2-1)-\log(x^3(y+1)^5)\]

OpenStudy (anonymous):

then turn the product in to a sum \[\frac{1}{2}\log(x^2-1)-\left(\log(x^3)+\log((y+1)^5)\right)\]

OpenStudy (anonymous):

factor and pull out the exponents as multipliers [\frac{1}{2}\left(\log((x+1)(x-1))-3\log(x)-5\log(x+1)\right)\]

OpenStudy (anonymous):

\[\frac{1}{2}\left(\log((x+1)(x-1))-3\log(x)-5\log(x+1)\right)\]

OpenStudy (anonymous):

one last step \[\frac{1}{2}\left(\log((x+1)+\log(x-1)-3\log(x)-5\log(x+1)\right)\]